OpenStudy (anonymous):

Perform the indicated operation. (8-15i)+(-3+2i)

Directrix (directrix):

(8-15i) + (-3+2i) = 8 - 15i - 3 + 2i = 8 - 3 - 15i + 2i = 5 - 13i which in "pure" a+bi form is --> 5 + (-13)i @bmx4life